UMAX-backend for SANE - User’s Guide

Oliver Rauch <Oliver.Rauch@Wolfsburg.DE>

6th march 1999

Contents

1 User’s Manual 3

1.1 Frontend functions and options: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.1.1 Scan modes: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.1.2 Scan sources: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.1.3 Options: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2 Installation Manual 6 2.1 Compilation and Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.2 Configuration: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3 SCSI adapter tips: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.4 Files: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.5 Environment: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.5.1 SANE DEBUG UMAX values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

3 Supported scanners 8

4 General Informations 10 4.1 BUGS: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.2 Authors: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 4.3 See also: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

1 User’s Manual

1.1 Frontend functions and options:

This backend dynamically enabels the options, scan-modes, etc. that are supported by the scanner in dependence of the scanning-mode and other options.

1.1.1 Scan modes: The available scan modes depend on the scanner type!

Mode  Remark
Lineart  1 bit/pixel black/white mode
Halftone 1 bit/pixel dithered black/white mode
Grayscale 8 / 10 / 12 / 14 bits/pixel gray mode
Color  24 / 30 / 36 / 48 bits/pixel RGB mode

1.1.2 Scan sources:

• Flatbed

• Transparency adapter (UTA)

• Automatic document feeder (ADF) The UTA and the ADF are only selectable if they are available!

1.1.3 Options:

There are two kind of options: the standard options that are often used and the advanced options.

Option  Remark
10-bit mode  Use 10 (30) bits/pixel instead of 8 (24) bits/pixel,

only supported by Astra 1200S.

(NOT SUPPORTED FROM ANY FRONTEND IN THE MOMENT!)

12-bit mode  Use 12 (36) bits/pixel instead of 8 (24) bits/pixel,

only supported by Astra 1220S.

(NOT SUPPORTED FROM ANY FRONTEND IN THE MOMENT!)

Analog gamma correction For color/intensity enhancement, value-range: 1..2,

doesn’t reduce the number of used colors!

only supported by some scanners (e.g. Supervista S12).

Bind RGB values  Use same values for each color, only selectable if such options are available
Brightness  Define brightness for halftone-mode.
Contrast  Define contrast for halftone-mode.
Custom gamma table  Digital gamma correction, for color/intensity enhancement!

Side-effect: may reduce the number of used colors!

For 30 and 36 bit scanners the used gamma table uses 30/36 input bits, so the reduction of the used colors is not so big!

Highlight  Define radiance level for highlight, only supported by some scanners.
Lamp warmup  Enable extended lamp warmup, only supported by some scanners.
Negative scan  Inverts color-intensity: for scanning negatives.
Quality calibration  Do quality white calibration.
Shadow  Define radiance level for shadow, only supported by some scanners.
Threshold  Define threshold for lineart-mode.

On graphical frontends the advanced options may be disabled so the frontend-window keeps small. To enable the advanced options take a look at the frontend’s manual!

Advanced Option  Remark
Cal. exposure time  Define exposure time for calibration,

only supported by some scanners.

Cal. lamp density  Define lamp density for calibration,

only supported by some scanners.

Scan exposure time  Define exposure time for scan,

only supported by some scanners.

Scan lamp density  Define lamp density for scan,

only supported by some scanners.

Set exposure time  Enable selection of exposure time,

if not enabled, scanner uses default values.

Set lamp density  Enable selection of lamp density, if not

enabled, scanner automatically selects a value.

Set scan lamp density Enable selection of lamp density for scan, if not enabled

scanner uses the value that was used for calibration.

2 Installation Manual

2.1 Compilation and Installation

The umax-backend is created and installed automatically while compiling and installing SANE.

2.2 Configuration:

The configuration file for the UMAX-backend resides in

     /usr/local/etc/sane.d/umax.conf.

Its contents is a list of device names that correspond to UMAX scanners. Empty lines and lines starting with a hash mark (#) are ignored. Lines starting with the word scsi use variable device selection, for further information take a look at sane-scsi.

A sample configuration file is shown below:

     #scsi vendor model type  bus channel id lun
     scsi  UMAX   *     *     *   *       *  *
     /dev/scanner
     # this is a comment
     /dev/sge

The special device name must be a generic SCSI device or a symlink to such a device. To find out to which device your scanner is assigned and how you have to set the permissions of that device, have a look at sane-scsi or use

the tool find-scanner.

The file may include devices, that correspond to scanners by other vendors. The Umax-backend will ignore these devices.

2.3 SCSI adapter tips:

The UMAX-scanners do block the scsi-bus for a few seconds while scanning. It is not necessary to connect the scanner to its own SCSI-adapter. But if you need short response time for your SCSI-harddisk (e.g. if your computer is a file-server), I suggest you use an own SCSI-adapter for your UMAX-scanner.

The SCSI-adapters that are shipped with some Umax-scanners are very simple ones. I suggest not to use it. If you really want to try: the UDS-IS11 is a DTC3181E-card based on the NCR5380-chip. On linux you may have the chance to get it work with the recent g NCR5380-driver. The card does not have IRQ/DMA. Try the following option for kernel/module: dtc3181e=0x280,254

See the relevant hardware FAQs and HOWTOs for your platform for more information.

See also: sane-scsi

2.4 Files:

• The backend configuration file: /usr/local/etc/sane.d/umax.conf

• The static library implementing this backend:

/usr/local/lib/libsane-umax.a

• The shared library implementing this backend :

/usr/local/lib/libsane-umax.so (present on systems that support dynamic loading)

2.5 Environment:

• SANE DEBUG UMAX

• see sane-dll

• see sane-scsi

• see backends

• see frontends

If the library was compiled with debug support enabled, this environment variable controls the debug level for this backend. E.g., a value of 128 requests all debug output to be printed. Smaller levels reduce verbosity:

2.5.1 SANE DEBUG UMAX values

Number Remark
0  print important errors (printed each time)
1  print errors
2  print sense
3  print warnings
4  print scanner-inquiry
5  print informations
6  print less important informations
7  print called procedures
8  print reader-process messages
10  print called sane-init-routines
11  print called sane-procedures
12  print sane infos
13  print sane option-control messages

Example: export SANE DEBUG UMAX=8

3 Supported scanners

The umax-backend only supports SCSI-scanners, parallel- and USB-scanners are not supported!

Name  Vendor SCSI-ID  F/W Line- Half- Gray- Color  Bits  Maximum
Version art tone scale RGB  resolution

UMAX scanners:

Vista S6  UMAX Vista-S6  all  ok  –  ok  ok  1,8,24  300 x 600
Vista S6E  UMAX Vista-S6E  all  ok  –  ok  ok  1,8,24  300 x 600
Vista S6E  UMAX UMAX S-6E  all  ok  –  ok  ok  1,8,24  300 x 600
Vista S6E  UMAX UMAX S-6EG  all  ok  –  ok  ok  1,8,24  300 x 600
Vista S8  UMAX Vista-S8  ?  ?  ?  ?  ?  1,8,24  400 x 800
Supervista S12  UMAX UMAX S-12G  all  ok  –  ok  ok  1,8,24  600 x 1200
Supervista S12  UMAX UMAX S-12  all  ok  –  ok  ok  1,8,24  600 x 1200
Supervista S12  UMAX Supervista S-12  all  ok  –  ok  ok  1,8,24  600 x 1200
Astra 600S  UMAX Astra 600S  all  ok  –  ok  ok  1,8,24  300 x 600
Astra 610S  UMAX Astra 610S  all  ok  –  ok  ok  1,8,24  300 x 600
Astra 1200S  UMAX Astra 1200S  all  ok  –  ok  ok  1,8,24,30 600 x 1200
Astra 1220S  UMAX Astra 1220S  all  ok  –  ok  ok  1,8,24,36 600 x 1200
UC630  UMAX UC630  not 1.6  ok  ?  bad  bad  1,8,24  300 x 300
UC630  UMAX UC630  1.6  ok  ?  ok  ok  1,8,24  300 x 300
UG630  UMAX UG630  all  ok  ?  ok  –  1,8  300 x 300
UG80  UMAX UG80  all  ok  ?  ok  –  1,8  300 x 300
UC840  UMAX UC840  not 1.6  ok  ?  bad  bad  1,8,24  400 x 400
UC840  UMAX UC840  1.6  ok  ?  ok  ok  1,8,24  400 x 400
UC1200S  UMAX UC1200S  ?  bad bad  bad  bad  1,8,24  600 x 600
UC1200SE  UMAX UC1200S  ?  ?  ?  ?  ?  1,8,24  600 x 600
UC1260  UMAX UC1260  not 1.6  ?  ?  ?  ?  1,8,24  600 x 1200
UC1260  UMAX UC1260  1.6  ok  ?  ok  ok  1,8,24  600 x 1200
Mirage  UMAX Mirage D-16L  ?  ?  ?  ?  ?  1,8,24,30 700 x 1400

1400 x 2800

Mirage II UMAX Mirage II ? ? ? ? ? 1,8,24,36 700 x 1400

1400 x 2800

Mirage IIse  UMAX Mirage IIse  ?  ?  ?  ?  ?  1,8,24,36 700 x 1400
Vista T630  UMAX Vista-T630  some  ok  –  ok  ok  1,8,24  300 x 600
PageScan  UMAX PSD  all  ok  –  ok  ok  1,8,24  300 x 300
PowerLook  UMAX unknown
PowerLook II  UMAX PL-II  some  ok  –  ok  ok  1,8,24,36 600 x 1200
PowerLook III  UMAX Power Look III  ?  ?  ?  ?  ?  1,8,24,42 1200 x 2400
PowerLook 2000  UMAX Power Look 2000  ?  ?  ?  ?  ?  1,8,24,36 1200 x 2400
PowerLook 3000  UMAX Power Look 3000  ?  ?  ?  ?  ?  1,8,24,42 1220 x 3048

3048 x 3048

Gemini D-16 UMAX Gemini D-16 ? ok – ok ok 1,8,24 400 x 800

800 x 1600

Linotype Hell scanners:

Jade  LinoHell Office  all  ok  –  ok  ok  1,8,24  600 x 1200
Jade2  LinoHell Office2  all  ok  –  ok  ok  1,8,24,30 600 x 1200
Saphir  LinoHell unknown
Saphir2 LinoHell SAPHIR2 all ok – ok ok 1,8,24,36 600 x 1200
Saphir Ultra  LinoHell unknown
Saphir Ultra II  LinoHell unknown
Saphir HiRes  LinoHell unknown
Opal  LinoHell unknown
Opal Ultra  LinoHell unknown

Vobis/Highscreen scanners:

Scanboostar Premium LinoHell Office2 all ok – ok ok 1,8,24,30 600 x 1200

Escom scanners:

Image Scanner 256 UMAX UG80 all ok ? ok – 1,8 300 x 300

Nikon scanners:

AX-210  Nikon  AX-210  all(?)  ok  ?  ok  –  1,8,24  600 x 1200

If you own a UMAX-scanner other than the ones listed above, it may work with SANE. It depends on the informations the scanner returns to the umax-backend. If the data-block is large enough, the backend prints a warning and continues, but it is possible that not everything works fine. I suggest you hold one hand on the power-button of the scanner while you try the first scans!

9

4 General Informations

4.1 BUGS:

• resolutions greater than 600 dpi sometimes don’t work!

• calibration by driver does not work for Astra 6X0S and Vista S6E

4.2 Authors:

Oliver Rauch

4.3 See also:

sane, introduction, backends, frontends, sane-scsi, sane-dll